# OpenPros: A Large-Scale Dataset for Limited View Prostate Ultrasound Computed Tomography

### v1

@@ Abstract
Prostate cancer is one of the most common and lethal cancers among men, making its early detection critically important. Although ultrasound imaging offers greater accessibility and cost-effectiveness compared to MRI, traditional transrectal ultrasound (TRUS) methods suffer from low sensitivity, especially in detecting anteriorly located tumors. Ultrasound computed tomography (USCT) provides quantitative tissue characterization, but its clinical implementation faces significant challenges, particularly under anatomically constrained limited-angle acquisition conditions specific to prostate imaging. To address these unmet needs, we introduce OpenPros, the first large-scale benchmark dataset explicitly developed for limited-view prostate USCT. Our dataset includes over 280,000 paired samples of realistic 2D speed-of-sound (SOS) phantoms and corresponding ultrasound full-waveform data, generated from anatomically accurate 3D digital prostate models derived from real clinical MRI/CT scans and ex vivo ultrasound measurements, annotated by medical experts. Simulations are conducted under clinically realistic configurations using advanced finite-difference time-domain (FDTD) and Runge-Kutta acoustic wave solvers, both provided as open-source components. Through comprehensive baseline experiments, we demonstrate that state-of-the-art deep learning methods surpass traditional physics-based approaches in both inference efficiency and reconstruction accuracy. Nevertheless, current deep learning models still fall short of delivering clinically acceptable high-resolution images with sufficient accuracy. By publicly releasing OpenPros, we aim to encourage the development of advanced machine learning algorithms capable of bridging this performance gap and producing clinically usable, high-resolution, and highly accurate prostate ultrasound images. The dataset is publicly accessible at https://open-pros.github.io/.
@@
Figure 1: OpenPros dataset creation and benchmarking pipeline. Top panel: Starting from clinical MRI and CT scans, we employ expert annotations to generate detailed 3D anatomical segmentations. We then incorporate real ultrasound speed-of-sound (SOS) measurements from ex vivo prostate samples acquired using the QTscan platform. These are integrated into comprehensive 3D abdominal SOS models. Clinically relevant 2D slices are subsequently extracted from these models to simulate limited-angle ultrasound tomography scenarios. Bottom panel: The extracted 2D SOS maps form the ground truth for ultrasound simulations governed by the acoustic wave equation. The resulting simulated ultrasound data are organized into the OpenPros dataset. We utilize these data to train and benchmark physics-based and deep-learning inversion methods, facilitating the evaluation and development of rapid, clinically relevant SOS reconstruction methods under challenging limited-angle conditions.
@@ 1 Introduction
Prostate cancer is the second most common malignancy among men and is one of the leading causes of cancer-related deaths worldwide, and every eight men suffers from it radtke2020re ; tosoian2024development . Since the 5-year survival rate for prostate cancer patients significantly drops from nearly 100% to approximately 34% once the disease progresses from localized or regional stages to distant metastases PCaStats , early detection of aggressive prostate cancer is of vital importance. Medical imaging plays an essential role in this early detection. Among the available imaging modalities, multiparametric MRI (mpMRI) is currently recognized as the most advanced and accurate imaging tool for detecting and localizing clinically significant prostate cancer. However, the high cost and limited accessibility of mpMRI restrict its widespread adoption, particularly in rural or low-resource settings de2014accuracy ; kasivisvanathan2018mri .
@@
In contrast, ultrasound imaging is widely accessible, cost-effective, and capable of real-time imaging. Prostate ultrasound is typically performed transrectally, producing B-mode (brightness-mode) images. Although transrectal ultrasound (TRUS) is the clinical standard for routine prostate evaluations and biopsy guidance, it has a sensitivity of only 30%–50% for detecting clinically significant tumors and a specificity of 70%–80% beemsterboer1999changing ; chen2016utility . Studies have further shown that tumors located in the anterior or apical prostate regions are often undetectable with TRUS due to poor soft-tissue contrast and restricted acoustic windows, and TRUS cannot reliably distinguish malignant lesions from benign conditions such as chronic prostatitis marivcic2010transrectal .
@@
Ultrasound computed tomography (USCT) has emerged as a promising alternative, reconstructing quantitative tissue parameters such as speed-of-sound (SOS) and acoustic attenuation, which serve as potential biomarkers for malignancy wu2024multiparametric ; williams2021mp22 . However, the anatomical constraints of prostate imaging inherently limit the acquisition aperture, creating a challenging scenario referred to as the limited-angle condition. Unlike idealized imaging setups where transducers surround the entire imaging domain, prostate imaging is anatomically restricted to transrectal and transabdominal placements, resulting in sparse and angularly limited data. Traditional physics-based methods typically struggle under these conditions, demonstrating slow convergence, severe ill-posedness, and significant reconstruction artifacts wang2025mask ; gilboy2020dual . Developing robust USCT algorithms capable of accurately handling limited-angle data is thus critically needed for clinical prostate imaging.
@@
Furthermore, clinical translation of prostate USCT faces considerable barriers due to the complexity and specialization of current imaging systems. To date, only two USCT systems (SoftVue and QTscan) have received U.S. FDA approval, and both systems focus exclusively on full-angle breast imaging with custom hardware setups unsuited for prostate applications sandhu2015frequency ; malik2018quantitative . These existing systems operate at relatively low frequencies, rely on patient positioning incompatible with prostate imaging, and require hours for reconstruction. Thus, there is an urgent need for efficient, generalizable, and clinically adaptable prostate-specific USCT platforms. Crucially, this advancement depends on the availability of realistic, anatomically precise digital phantoms and datasets, which are currently lacking in the field gilboy2020dual ; aalamifar2017enabling .
@@
Recent advances in deep learning, particularly convolutional neural networks (CNNs), have shown potential for overcoming these limitations by learning complex mappings directly from ultrasound data to high-resolution SOS maps chugh2021survey ; havaei2017brain . Data-driven approaches bypass computational bottlenecks encountered by iterative solvers and demonstrate the ability to reconstruct detailed tissue properties even under sparse and noisy acquisition conditions. Instead of hours to days of image reconstruction using physics-based methods and the requirement of expert reading as a follow-up, the relatively short inference time and the automatic analysis enable faster and easier diagnosis for better patient experience. Furthermore, transformer-based architectures have recently demonstrated remarkable performance in medical imaging by effectively modeling long-range spatial dependencies, a feature particularly beneficial for ultrasound tomography due to the extensive spatial interaction of acoustic waves.
Despite these advancements, progress has been significantly hampered by the lack of large-scale, high-fidelity datasets supporting the development, evaluation, and reproducibility of innovative reconstruction algorithms. Existing USCT datasets are typically for breast imaging, which are either synthesized in simulation or derived from real phantoms. Related datasets are listed in Table 1. They exhibit diversities and reflect anatomical realism to a good extent but not the best. There are also anatomical datasets for the male pelvic region but not for USCT purposes. The commercial anatomy softwares such as Zygote Body and Complete Anatomy provide different pricing options for viewing and downloading, but they generally lack anatomical varieties. Last but not least, no publicly available dataset adequately addresses the unique challenges posed by limited-view prostate ultrasound computed tomography and the existence of bones in the imaging view while simultaneously providing realistic wave propagation modeling and comprehensive full-waveform data.
@@
Dataset Prostate Acoustic parameters Actual anatomy Tissue heterogeneity Bones Limited angle Public Free access OpenPros (ours) Li et al. li20213 Ruiter et al. ruiter2018usct OpenWaves zeng2025openwaves Segars et al. segars20104d NA NA The visible human project ackerman1998visible NA NA Zygote Body NA NA Complete Anatomy NA NA
Motivated by these challenges and inspired by the success of OpenFWI deng2022openfwi and $\mathbf{E}^{FWI}$ feng2023mathbf datasets, we introduce OpenPros, the first large-scale dataset specifically designed for limited-angle prostate USCT scenarios. A schematic illustration of the overall pipelines of OpenPros is shown in Figure 1. Our dataset comprises over 280,000 paired 2D SOS phantoms and ultrasound full-waveform data derived from anatomically realistic 3D prostate models generated from clinical MRI/CT scans and ex vivo ultrasound measurements, annotated meticulously by clinical experts. OpenPros serves as a critical benchmark facilitating advances in computational efficiency, limited-angle reconstruction accuracy, rapid clinical adaptability, and comprehensive method comparisons across various imaging conditions, including ray-based, single scattering, high-frequency, and limited-angle scenarios.
@@
    Comprehensive benchmarking of inversion methods: Thorough evaluation of physics-based and state-of-the-art deep learning methods under realistic limited-angle conditions, establishing clear performance baselines and guiding future algorithmic improvements.
@@ 2 Ultrasound Computed Tomography and Forward Modeling
$\nabla^{2}p-\frac{1}{c^{2}}\frac{\partial^{2}p}{\partial t^{2}}=s,$ (1)
@@
Data-driven USCT leverages neural networks to directly approximate the inverse mapping $c=f^{-1}(p)$, as demonstrated in recent studies wu2019inversionnet . Thus, the specific task addressed in this paper is the supervised learning problem, formulated as $\min_{\theta}\mathbb{E}{p,c}\left[\mathcal{L}(c,\hat{c})\right]$, where $\hat{c}=f{\theta}^{-1}(p)$. Here, $f_{\theta}^{-1}$ represents a deep neural network parameterized by $\theta$, trained on pairs of simulated ultrasound signals $p$ and corresponding ground-truth SOS maps $c$. The training objective $\mathcal{L}$ typically incorporates quantitative metrics such as Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), Structural Similarity Index Measure (SSIM) and Pearson Correlation Coefficient (PCC).
@@ 3 OpenPros Dataset
OpenPros is the first large-scale benchmark dataset explicitly designed to facilitate research in limited-angle prostate ultrasound computed tomography (USCT). It contains anatomically realistic 2D speed-of-sound phantoms, organ segmentation labels, and corresponding simulated ultrasound waveforms derived from detailed 3D digital prostate models. The patient level anatomy ID is named as 3_01, 3_02, 3_03 and 3_04. The prostate level anatomy ID is named as the date of acquisition (in total 62 prostates). Detailed naming strategy an be found in the Appendix C.
@@ 3.1 Dataset Statistics
 OpenPros 6.8 TB 224K / 28K / 28K $1140\times 40\times 1000\times 161$ $1140\times 1\times 401\times 161$
@@
Anatomy-Aware Inversion: Integrated segmentation masks guide targeted inversion studies, emphasizing reconstruction accuracy in clinically significant regions, thus promoting anatomically detailed reconstruction assessments.
@@
It is important to note that the 3D digital phantoms were derived from human CT/MRI nyholm2018mr and USCT scans of ex vivo prostate specimens parikh2024mp30 ; wiskin2022imaging ; williams2021mp22 . Major organs were annotated by experts using T2-weighted MRI, fat was segmented from T1-weighted MRI, bones were segmented from X-ray CT, and the speed of sound and attenuation of ex vivo prostate samples were measured using a QT scanner (QT Imaging Inc., Novato, California, USA). Speed of sound of other organs were acquired from ITIS foundation tissue database baumgartner2024database . To best mimic the tissue heterogeneity, we employed Gaussian distributions with given mean values and standard deviations from the tissue database to assign speed of sound in different tissue types. These derivations from human data reduces the reliance on synthetic simulation and maximizes the fidelity of the dataset, especially in the prostate area. Additional details on phantom construction, applications, and open access availability can be found in wu2024realistic . The 3D abdominal anatomical structure and the ultrasound probe placement sketch can be found in Figure 2(a) and the illustration of 3D phantom can be found in Figure 2(b).
@@
We generate 2D speed-of-sound phantoms by slicing our anatomically realistic 3D prostate volumes under clinically realistic probe configurations. For each phantom, we place one transrectal probe in the rectum and one abdominal probe on the body surface, then sample cross-sections across a range of rotations (±$\pm 45^{\circ}$) and small random perturbations. This process yields 280 K paired 2D SOS maps and corresponding ultrasound waveforms that faithfully capture patient-specific anatomy and limited-angle acquisition variability.
More detailed 2D phantom extraction strategy can be found in Appendix B
@@ 3.4 Ultrasound Data Simulation
Figure 3: Examples of simulated ultrasound data and phantoms: without (top) and with (bottom) bone in the phantoms. We show two example channels of reflections and transmissions with sources (yellow stars) and receivers (red dots) on the two probes. Our PDE solvers can simulate complex and realistic ultrasound wave phenomena, such like transmissions, reflections, direct waves and multi-scatterings.
The spatial discretization of the domain uses a grid spacing of 0.375 mm, resulting in a field of view of approximately 60 mm in depth and 150 mm in width. This configuration mirrors the anatomical scale of the prostate and its surrounding structures. A summary of the physical simulation parameters and the physical meaning of the dimension is shown in Table 3.
@@
 Dataset Grid Spacing SOS Map Spatial Size Source Spacing Source Line Length Receiver Spacing Receiver Line Length Time Spacing Recorded Time  OpenPros 0.375 mm 60 mm $\times$ 150 mm 37.5 mm 60 mm 3.75 mm 60 mm $1\times 10^{-7}$ s $100~{}\mu s$

### v2

@@ Abstract
Prostate cancer is one of the most common and lethal cancers among men, making its early detection critically important. Ultrasound computed tomography (USCT) has emerged as an accessible and cost-effective method that reconstructs quantitative tissue parameters, which can serve as potential biomarkers for malignancy. However, current prostate USCT faces considerable barriers: limited-angle acquisitions due to anatomical constraints, tissue heterogeneity, proximity to organs and bony pelvic structures, and lengthy processing times. The lack of large-scale, anatomically precise datasets significantly hampers the development of high-quality, efficient, and generalizable methods. To address this gap, we introduce OpenPros, the first large-scale benchmark dataset for limited-angle prostate USCT, designed to evaluate machine learning algorithms for inverse problems systematically. Our dataset includes over 280,000 paired samples of realistic 2D speed-of-sound (SOS) phantoms and corresponding ultrasound full-waveform data, generated from anatomically accurate 3D digital prostate models derived from 4 real clinical MRI/CT scans and 62 ex vivo prostate specimens with experimental ultrasound measurements, annotated by medical experts. Simulations are conducted under clinically realistic configurations using advanced finite-difference time-domain (FDTD) and Runge-Kutta acoustic wave solvers, both provided as open-source components. Through comprehensive benchmarking, we find that deep learning methods significantly outperform traditional physics-based algorithms in inference efficiency and reconstruction accuracy. However, our results also reveal that current machine learning methods fail to deliver clinically acceptable, high-resolution reconstructions, underscoring critical gaps in generalization, robustness, and uncertainty quantification. By publicly releasing OpenPros, we provide the community with a rigorous benchmark that not only enables fair method comparison but also motivates new advances in physics-informed learning, foundation models for scientific imaging, and uncertainty-aware reconstruction—bridging the gap between academic ML research and real-world clinical deployment. The dataset is publicly accessible at https://open-pros.github.io/.
@@
Figure 1: OpenPros dataset creation and benchmarking pipeline. Top panel: Starting from clinical MRI and CT scans, we employ expert annotations to generate detailed 3D anatomical segmentations. We then incorporate real ultrasound speed-of-sound (SOS) measurements from ex vivo prostate samples acquired using the QTscan platform. These are integrated into comprehensive 3D abdominal SOS models. Clinically relevant 2D slices are extracted from these models to simulate limited-angle ultrasound tomography scenarios. Bottom panel: The extracted 2D SOS maps form the ground truth for ultrasound simulations governed by the acoustic wave equation. The resulting simulated ultrasound data are organized into the OpenPros dataset. We utilize these data to train and benchmark physics-based and deep-learning inversion methods, facilitating the evaluation and development of rapid, clinically relevant SOS reconstruction methods under challenging limited-angle conditions.
@@ 1 Introduction
Prostate cancer is the second most common malignancy in men and is one of the leading causes of cancer-related deaths worldwide. One in eight men suffers from it (Radtke and Hadaschik, 2020; Tosoian et al., 2024). Since the 5-year survival rate for prostate cancer patients significantly drops from nearly 100% to approximately 34% once the disease progresses from localized or regional stages to distant metastases (Institute, 2024), early detection of aggressive prostate cancer is of vital importance. Medical imaging plays an essential role in this early detection. Among the available imaging modalities, multiparametric MRI (mpMRI) is currently recognized as the most advanced and accurate imaging tool for detecting and localizing clinically significant prostate cancer. However, the high cost and limited accessibility of mpMRI restrict its widespread adoption, particularly in rural or low-resource settings (De Rooij et al., 2014; Kasivisvanathan et al., 2018).
@@
In contrast, ultrasound imaging is widely accessible, cost-effective, and capable of real-time imaging. Prostate ultrasound is typically performed transrectally, producing B-mode (brightness-mode) images. Although transrectal ultrasound (TRUS) is the clinical standard for routine prostate evaluations and biopsy guidance, it has a sensitivity of only 30%–50% for detecting clinically significant tumors and a specificity of 70%–80% (Beemsterboer et al., 1999; Chen et al., 2016). Studies have further shown that tumors located in the anterior or apical prostate regions are often undetectable with TRUS due to poor soft-tissue contrast and restricted acoustic windows, and TRUS cannot reliably distinguish malignant lesions from benign conditions such as chronic prostatitis (Maričić et al., 2010).
@@
Ultrasound computed tomography (USCT) has emerged as a promising alternative, reconstructing quantitative tissue parameters like speed-of-sound (SOS) and acoustic attenuation that serve as potential biomarkers for malignancy (Wu, 2024; Williams et al., 2021). However, the anatomical constraints of prostate imaging inherently limit the acquisition aperture, creating a challenging limited-angle condition. Unlike idealized setups where transducers surround the entire imaging domain, prostate imaging is anatomically restricted to transrectal and transabdominal placements, resulting in sparse and angularly limited data. Traditional physics-based methods typically struggle under these conditions, with slow convergence, severe ill-posedness, and significant reconstruction artifacts (Wang et al., 2025; Gilboy et al., 2020). Developing robust USCT algorithms capable of accurately handling limited-angle data is thus critically needed for clinical prostate imaging.
@@
Furthermore, clinical translation of prostate USCT faces considerable barriers due to the complexity and specialization of current imaging systems. To date, only two USCT systems (SoftVue and QTscan) have received U.S. FDA approval, and both systems focus exclusively on full-angle breast imaging with custom hardware setups unsuited for prostate applications (Sandhu et al., 2015; Malik et al., 2018). These existing systems operate at relatively low frequencies, rely on patient positioning incompatible with prostate imaging, and require hours for reconstruction. Thus, there is an urgent need for efficient, generalizable, and clinically adaptable prostate-specific USCT platforms. Crucially, this advancement depends on the availability of realistic, anatomically precise digital phantoms and datasets, which are currently lacking in the field (Gilboy et al., 2020; Aalamifar et al., 2017).
@@
Recent advances in deep learning, particularly convolutional neural networks (CNNs), have shown potential for overcoming these limitations by learning complex mappings directly from ultrasound data to high-resolution SOS maps (Chugh et al., 2021; Havaei et al., 2017). Data-driven approaches bypass computational bottlenecks encountered by iterative solvers and demonstrate the ability to reconstruct detailed tissue properties even under sparse and noisy acquisition conditions. Instead of hours to days of image reconstruction using physics-based methods and the requirement of expert reading as a follow-up, the relatively short inference time and the automatic analysis enable faster and easier diagnosis for better patient experience. Furthermore, transformer-based architectures have recently demonstrated remarkable performance in medical imaging by effectively modeling long-range spatial dependencies, a feature particularly beneficial for ultrasound tomography due to the extensive spatial interaction of acoustic waves.
Despite these advancements, progress has been significantly hampered by the lack of large-scale, high-fidelity datasets supporting the development, evaluation, and reproducibility of innovative reconstruction algorithms. Existing USCT datasets are typically for breast imaging, which are either synthesized in simulation or derived from real phantoms. Related datasets are listed in Table 1. They exhibit diversity and reflect anatomical realism to a good extent, they are not optimal for developing and benchmarking advancing prostate USCT algorithm. There are also anatomical datasets for the male pelvic region but not for USCT purposes. The commercial anatomy softwares such as Zygote Body and Complete Anatomy provide different pricing options for viewing and downloading, but they generally lack anatomical varieties. Moreover, no publicly available dataset adequately addresses the unique challenges posed by limited-view prostate USCT and the existence of bones in the imaging view while simultaneously providing realistic wave propagation modeling and comprehensive full-waveform data.
@@
Dataset Prostate Acoustic parameters Actual anatomy Tissue heterogeneity Bones Limited angle Public Free access OpenPros (ours) Li et al. Li et al. (2021) Ruiter et al. (Ruiter et al., 2018) OpenWaves (Zeng et al., 2025) Segars et al. (Segars et al., 2010) NA NA The visible human project (Ackerman, 1998) NA NA Zygote Body NA NA Complete Anatomy NA NA
Motivated by these challenges and the pressing need for prostate USCT, we introduce OpenPros, the first large-scale dataset specifically designed for limited-angle prostate USCT scenarios. A schematic illustration of the overall pipelines of OpenPros is shown in Figure 1. Our dataset comprises over 280,000 paired 2D SOS phantoms and ultrasound full-waveform data derived from anatomically realistic 3D prostate models generated from 4 clinical MRI/CT scans and 62 ex vivo prostate specimens with experimentally measured speed-of-sound (SOS), annotated meticulously by clinical experts. OpenPros serves as a critical benchmark facilitating advances in computational efficiency, limited-angle reconstruction accuracy, rapid clinical adaptability, and comprehensive method comparisons across various imaging conditions, including ray-based, single scattering, high-frequency, and limited-angle scenarios.
@@
    Comprehensive benchmarking of inversion methods: Thorough evaluation of physics-based and deep learning methods under realistic limited-angle conditions, including systematic tests of generalization, robustness, and inference efficiency. These baselines establish clear performance baselines and guide future algorithmic improvements.
@@ 2 Ultrasound Computed Tomography and Forward Modeling
$\vskip-2.84526pt\nabla^{2}p-\frac{1}{c^{2}}\frac{\partial^{2}p}{\partial t^{2}}=s,$ (1)
@@
[…]
@@ 3 OpenPros Dataset
OpenPros is the first large-scale benchmark dataset explicitly designed to facilitate research in limited-angle prostate ultrasound computed tomography (USCT). It contains anatomically realistic 2D speed-of-sound phantoms, organ segmentation labels, and corresponding simulated ultrasound waveforms derived from detailed 3D digital prostate models. The patient level anatomy ID is named as 3_01, 3_02, 3_03 and 3_04. The prostate level anatomy ID is named as the date of acquisition (in total 62 prostates). Detailed naming strategy can be found in the Appendix A.2. Example data pairs and FDTD simulation code are provided in the supplementary materials
@@ 3.1 Dataset Statistics
 OpenPros 6.8 TB 224K / 28K / 28K $\times 40\times 1000\times 161 1140×1×401$$\times 1\times 401\times 161$
@@
It is important to note that the 3D digital phantoms were derived from human CT/MRI (Nyholm et al., 2018) and USCT scans of ex vivo prostate specimens (Parikh et al., 2024; Wiskin et al., 2022; Williams et al., 2021). Major organs were annotated by experts using T2-weighted MRI, fat was segmented from T1-weighted MRI, bones were segmented from X-ray CT, and the speed of sound and attenuation of ex vivo prostate samples were measured using a QT scanner (QT Imaging Inc., Novato, California, USA). Speed of sound of other organs were acquired from ITIS foundation tissue database (Baumgartner et al., 2024). To best mimic the tissue heterogeneity, we employed Gaussian distributions with given mean values and standard deviations from the tissue database to assign speed of sound in different tissue types. These derivations from human data reduces the reliance on synthetic simulation and maximizes the fidelity of the dataset, especially in the prostate area. Additional details on phantom construction, applications, and open access availability can be found in (Wu et al., 2024). The 3D abdominal anatomical structure and the ultrasound probe placement sketch can be found in Figure 2(a) and the illustration of 3D phantom can be found in Figure 2(b).
@@
We generate 2D speed-of-sound phantoms by slicing our anatomically realistic 3D prostate volumes under clinically realistic probe configurations. For each phantom, we place one transrectal probe in the rectum and one abdominal probe on the body surface, then sample cross-sections across a range of rotations ($\pm 45^{\circ}$) and small random perturbations. This process yields 280 K paired 2D SOS maps and corresponding ultrasound waveforms that faithfully capture patient-specific anatomy and limited-angle acquisition variability.
More detailed 2D phantom extraction strategy can be found in Appendix A.1
@@ 3.4 Ultrasound Data Simulation
Figure 3: Examples of simulated ultrasound data and phantoms: without (top) and with (bottom) bone in the phantoms. We show two example channels of reflections and transmissions with sources (yellow stars) and receivers (red dots) on the two probes. Our PDE solvers can simulate complex and realistic ultrasound wave phenomena, including transmissions, reflections, direct waves, and multi-scatterings.
The spatial discretization of the domain uses a grid spacing of 0.375 mm, resulting in a field of view of approximately 60 mm in width and 150 mm in depth. Each SOS map has a spatial resolution of 401 × 161 grid points, corresponding to a physical field of view of 60 mm (lateral) by 150 mm (axial), with uniform grid spacing of 0.375 mm in the lateral and axial directions. These physical dimensions and spacings are kept constant across all examples in the dataset. This configuration mirrors the anatomical scale of the prostate and its surrounding structures. A summary of the physical simulation parameters and the physical meaning of the dimension is shown in Table 3.
@@
 Dataset Grid Spacing SOS Map Spatial Size Source Spacing Source Line Length Receiver Spacing Receiver Line Length Time Spacing Recorded Time  OpenPros 0.375 mm 60 mm $\times$ 150 mm 3.75 mm 60 mm 0.375 mm 60 mm $1\times 10^{-7}$ s $100~\mu s$